OpenStudy (anonymous):

what is the square root of 125?

OpenStudy (anonymous):

What do you think it is? Remember that this is not a perfect square.

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

\[=\sqrt{25*5}=5\sqrt{5}\]

OpenStudy (anonymous):

@Raia

OpenStudy (anonymous):

Oh i was gone for a minute

OpenStudy (anonymous):

i found the answer, thanks